# On the Verification of Control Flow Attestation Evidence

### v1

@@
Abstract
Remote run-time attestation methods, including Control Flow Attestation (CFA) and Data Flow Attestation (DFA), have been proposed to generate precise evidence of execution’s control flow path (in CFA) and optionally execution data inputs (in DFA) on a remote and potentially compromised embedded device, hereby referred to as a Prover ($\sf{\mathcal{P}rv}$). Recent advances in run-time attestation architectures are also able to guarantee that a remote Verifier ($\sf{\mathcal{V}rf}$) reliably receives this evidence from $\sf{\mathcal{P}rv}$, even when $\sf{\mathcal{P}rv}$’s software state is fully compromised. This, in theory, enables secure “run-time auditing” in addition to best-effort attestation, i.e., it guarantees that $\sf{\mathcal{V}rf}$ can examine execution evidence to identify previously unknown compromises as soon as they are exploited, pinpoint their root cause(s), and remediate them. However, prior work has for the most part focused on securely implementing $\sf{\mathcal{P}rv}$’s root of trust (responsible for generating authentic run-time evidence), leaving $\sf{\mathcal{V}rf}$’s perspective in this security service unexplored.
@@ Abstract
In this work, we argue that run-time attestation and auditing are only truly useful if $\sf{\mathcal{V}rf}$ can effectively analyze received evidence. From this premise, we characterize different types of evidence produced by existing run-time attestation/auditing architectures in terms of $\sf{\mathcal{V}rf}$’s ability to detect and remediate (previously unknown) vulnerabilities. As a case study for practical uses of run-time evidence by $\sf{\mathcal{V}rf}$, we propose SABRE: a Security Analysis and Binary Repair Engine. SABRE showcases how $\sf{\mathcal{V}rf}$ can systematically leverage run-time evidence to detect control flow attacks, pinpoint corrupted control data and specific instructions used to corrupt them, and leverage this evidence to automatically generate binary patches to buffer overflow and use-after-free vulnerabilities without source code knowledge.
Index Terms:
Control Flow Attestation, Systems Security, Embedded Systems
1 Introduction
@@
Embedded devices exist at the edge of larger systems, often serving as remotely operated sensors and on-demand actuators. Due to low energy and cost requirements, they are commonly implemented using low-end micro-controller units (MCUs) that lack the architectural security features of application computers (e.g., virtual memory or memory management units). Consequently, MCUs are relatively more vulnerable and frequently targeted by cyber attacks [1, 2].
@@
Given their low budget for preventative security features, MCUs need inexpensive means to prove their software integrity to back-ends that rely on their services (e.g., a device owner or control center). Towards this goal, Remote Attestation (RA[3] has been proposed to give a resource-rich back-end – called a Verifier ($\sf{\mathcal{V}rf}$) – the ability to remotely assess the software state of a resource-constrained Prover MCU ($\sf{\mathcal{P}rv}$). In RA, $\sf{\mathcal{V}rf}$ challenges $\sf{\mathcal{P}rv}$ to return a cryptographic proof of $\sf{\mathcal{P}rv}$’s currently installed software binary. Only after receiving an authentic proof will $\sf{\mathcal{V}rf}$ trust that $\sf{\mathcal{P}rv}$ is installed with the correct software.
@@
Although RA can effectively detect illegal code modifications, it cannot detect run-time attacks that do not modify code [4]. For example, an adversary ($\sf{\mathcal{A}dv}$) could hijack the control flow of a program by exploiting a memory vulnerability (e.g., a buffer overflow [5]) to corrupt control data, including return addresses, indirect jump targets, and function pointers. Consequently, $\sf{\mathcal{A}dv}$ can skip access controls checks [6] or chain the out-of-order execution of sub-sequences of instructions to implement (often Turing-complete) malicious behavior [7, 8]. Since these attacks do not modify code, they are oblivious to RA.
@@
To address this limitation, RA can be extended into Control Flow Attestation (CFA). In CFA, $\sf{\mathcal{V}rf}$ obtains cryptographic proof of both the installed software image and the exact control flow path followed in the most recent execution of the software [9]. This evidence and the respective proof are generated/authenticated by a root of trust (RoT) within $\sf{\mathcal{P}rv}$ that securely records control flow transfer destinations. Therefore, CFA evidence can, in principle, allow $\sf{\mathcal{V}rf}$ to detect both code modification and out-of-order execution attacks. Recent architectures augment CFA to obtain “Run-time Auditing” [6, 10] by guaranteeing the delivery of CFA evidence to $\sf{\mathcal{V}rf}$ even when $\sf{\mathcal{P}rv}$’s software is compromised. See Sec. 2 for details on CFA and run-time auditing RoTs.
We note, however, that prior work in this area [9] has primarily focused on secure design/implementation of the $\sf{\mathcal{P}rv}$ RoTs, often overlooking $\sf{\mathcal{V}rf}$’s role in run-time attestation/auditing. Current efforts focusing on $\sf{\mathcal{V}rf}$ designs look into hiding control flow paths while enabling simple remote checks via Zero-Knowledge proofs [11] or addressing scalability concerns of building control flow graphs (CFG) for path verification [12].
@@
Yet, despite the promising benefits of CFA, no prior work has systematically analyzed $\sf{\mathcal{V}rf}$ ability to interpret received CFA evidence to identify root cause memory vulnerabilities that lead to attacks and remediate them. These challenges are catalyzed when firmware source codes are unavailable due to containing proprietary dependencies/third-party libraries: an unfortunate common occurrence in embedded system software chains. Remarkably, given the prevalence of memory-unsafe languages in embedded implementations and their various associated exploits and bugs [13], we argue that automated methods to process run-time evidence to identify unknown memory safety vulnerabilities (as soon as they are exploited in control flow attacks) can result in timely detection of otherwise oblivious attacks and elimination of their causes.
While techniques have been proposed for root cause vulnerability analysis in high-end devices (e.g., based on Intel PT traces [14, 15]) there are no frameworks for analysis of execution traces produced by run-time attestation in MCUs. Our work aims to bridge these gaps with the following intended contributions:
@@
    We analyze and classify the types of run-time evidence generated by existing CFA RoTs. Their trade-offs are examined in detail with particular focus on their ability to support $\sf{\mathcal{V}rf}$’s remote detection and remediation of vulnerabilities.
@@
    As a case study, we present SABRE: a Security Analysis and Binary Repair Engine based on CFA evidence. SABRE combines binary analysis with CFA run-time evidence to automatically locate and patch buffer overflow and use-after-free vulnerabilities without requiring source-code knowledge. SABRE analysis combines backward definitions analysis along with forward symbolic data-flow analysis to identify such root causes. Based on this analysis and the detected vulnerability, SABRE automatically generates patches and validates their effectiveness.
@@
    We develop a fully functional and open-source prototype of SABRE with support for binaries of TI MSP430 and ARM Cortex-M MCU families [16]. SABRE prototype is also directly compatible with run-time evidence generated by existing open-source run-time attestation RoTs based on Trusted Execution Environments (TEEs) [10] and custom hardware [6].
2 Background
2.1 Remote Attestation
@@
    Upon receiving $Chal$, an RoT in $\sf{\mathcal{P}rv}$ computes an authenticated integrity-ensuring function over its program memory and $Chal$ to produce report $H$.
@@
A RoT in $\sf{\mathcal{P}rv}$ performs the authenticated measurement (step 2) by generating an integrity token, e.g., a message authentication code (MAC) or a digital signature, over program memory using a secret key (either pre-shared symmetric key or the private key for a public key $\sf{\mathcal{V}rf}$ knows). Since threat models of RA (and of CFA consequently) assume that $\sf{\mathcal{P}rv}$ is susceptible to full software compromise, the secret key used in this operation must be securely stored by $\sf{\mathcal{P}rv}$’s RoT. Secure storage for the RA secret key implies some level of hardware support for the RoT implementation, ensuring that the key is unmodifiable and inaccessible to any untrusted software running on $\sf{\mathcal{P}rv}$.
Some early RA proposals aimed to generate evidence of $\sf{\mathcal{P}rv}$’s code integrity without hardware support. In this setting, the evidence provided to $\sf{\mathcal{V}rf}$ must be deterministic and have a measurable architectural side-effect (e.g., predictable delivery time [17]). These “software-based” architectures detect attempts to forge RA evidence based on delayed responses from $\sf{\mathcal{P}rv}$. Hence, they require assumptions about deterministic timing characteristics, direct wired interfaces between $\sf{\mathcal{V}rf}$-$\sf{\mathcal{P}rv}$, and predictable latency [17, 18, 19, 20]. Unfortunately, these strict requirements make them incompatible with remote settings [21]. Recent software-based architectures replace the timing requirements for a trusted software boot-strapping phase [22, 23]; additionally, they use software-fault isolation to ensure the integrity of the trusted components at run-time.
Figure 1: RA interaction
Hardware-based RA [24, 25, 26, 27, 28] incorporates dedicated hardware to securely create RA evidence that depends on secrets rather than timing characteristics. This is achieved through external modules (such as custom hardware extensions [27, 28, 25] or TPMs [26]) or RA-dedicated co-processors [24]. While they provide stronger security guarantees than earlier software-based techniques, the added hardware cost can be prohibitive for resource-constrained devices, such as MCUs.
Hybrid RA schemes  [29, 30, 31] use a combination of minimal hardware components and trusted software modules to achieve the security guarantees equivalent to hardware-based RA at a manageable hardware cost. Hybrid approaches typically implement the RA measurement (e.g., MAC or signature functionality) in software, while a low-cost hardware monitor protects the integrity of the RA measurement code and its execution and guarantees the secrecy of the RA cryptographic key(s).
2.2 Run-Time Code Reuse/Control Flow Attacks
[…]
2.2.1 Control Flow Attack Root Causes and their Prevalence in Embedded Systems Software
One class of memory vulnerabilities that can be exploited to launch a control flow attack is a buffer overflow [5]. A buffer overflow occurs when a memory write exceeds the bounds of a buffer. If $\sf{\mathcal{A}dv}$ can cause a buffer overflow and control its input, they can leverage this vulnerability to overwrite adjacent memory that stores indirect branch destinations (e.g., return addresses or function pointers), successfully launching control flow attacks without injecting code directly. See Appendix A.1 for a concrete example of a control flow attack via a buffer overflow.
Buffer overflows are one of the most commonly exploited memory vulnerabilities [36], especially in embedded systems software that is commonly written in unsafe memory languages like C/C++. Some recent examples include an exploited buffer overflow in Realtek chipset’s SDK that allowed remote attackers to launch arbitrary code execution in routers and various IoT devices [37] (consequently resulting in CVE-2021-35395 [38]). Buffer overflow vulnerabilities were also previously disclosed in embedded real-time operating systems (RTOS), leading to arbitrary code execution (e.g., CVE-2020-10023 [39] and CVE-2020-10019 [40] in Zephyr [41]). This style of attack is not limited to return addresses on the stack. For example, each instance of a C++ object includes a function pointer to a “vtable” of the object’s C++ class’s functions [42]. Thus, an $\sf{\mathcal{A}dv}$ could hijack a program’s control flow by causing an overflow into a C++ object’s function pointers. This applies equally to C structs containing function pointers.
Another memory vulnerability that often leads to control flow attacks is use-after-free [43]. In use-after-free, dynamically allocated memory is used after it has been freed, typically leading to a program crash when accidental (i.e., in a non-malicious case). When exploited, it can corrupt control data or launch control flow attacks. For example, when that freed memory previously contained function/vtable pointers (e.g., in the case of dynamically allocated C structs or C++ objects), $\sf{\mathcal{A}dv}$ can modify freed memory so that the reuse of this previously freed memory leads to a control flow hijack. A concrete example of how a use-after-free can lead to control flow hijacking is presented in Appendix A.2.
Use-after-frees were ranked as one of the most dangerous software weaknesses by MITRE [36]. Embedded libraries/firmware that use dynamic allocation [44, 45, 46] can be equally vulnerable use-after-free vulnerabilities [47, 48]. An example of use-after-free leading to arbitrary code execution in embedded platforms includes CVE-2017-14201 [47] that affected Zephyr RTOS [41].
Whether launched by a buffer overflow or a use-after-free vulnerability, such attacks do not require code modifications. Thus, they would remain undetected by classic RA mechanisms.
2.3 Run-time Attestation and Auditing
@@
Run-time Attestation and Auditing are a set of security services that aim to generate and guarantee the delivery of accurate and authentic evidence of a $\sf{\mathcal{P}rv}$’s run-time behavior to a remote $\sf{\mathcal{V}rf}$, thus reflecting to $\sf{\mathcal{V}rf}$ whether any control flow attack has occurred. One type of run-time attestation is Control Flow Attestation (CFA[49, 50, 51, 52, 53, 54, 4, 55, 56, 57, 58], which extends RA to detect run-time attacks in addition to illegal binary modifications. Upon receiving a request to execute some software function along with a RA challenge from $\sf{\mathcal{V}rf}$, $\sf{\mathcal{P}rv}$ executes the requested function, and a RoT in $\sf{\mathcal{P}rv}$ records a trace of this execution into protected memory. In the most fine-grained case, the trace is a control flow log ($\mathcal{C}F_{Log}$) of all control flow transfers that have occurred during the requested execution. After producing $\mathcal{C}F_{Log}$, $\sf{\mathcal{P}rv}$ performs the authenticated RA integrity measurement over the $\mathcal{C}F_{Log}$, the received challenge, and its own program memory to produce a timely and authenticated report. This report attests to the static state of the $\sf{\mathcal{P}rv}$’s program binary and the run-time state within this timely execution. When applicable, data inputs – in data flow attestation (DFA[50, 59, 51] – or produced outputs [60, 61] can also be included in this proof to bind obtained results to the proper execution of respective software functions on expected inputs.
@@
Existing CFA techniques use either (1) binary instrumentation along with TEE support; or (2) custom hardware modifications to generate $\mathcal{C}F_{Log}$ by detecting and saving each branch destination to a dedicated and protected memory region. For techniques that use binary instrumentation, a pre-processing phase modifies the binary so that all branch instructions (e.g., jumps, returns, calls, etc.) are prepended with additional calls to a TEE-protected trusted code. Once called, the trusted code appends $\mathcal{C}F_{Log}$ to update the trace with the current branch destination. In hardware-based techniques, custom hardware interfaces with the MCU core to detect branches and records their destinations to a reserved memory region. Recent work [6, 10] has also created mechanisms to implement reliable communication with $\sf{\mathcal{V}rf}$ as a part of the CFA RoT, enabling “run-time auditing” by ensuring evidence delivery even in the presence of malware that infects $\sf{\mathcal{P}rv}$ and refuses to deliver evidence about its actions/presence.
3 Characterizing Run-time Evidence and their Utility
3.1 Evidence Types

### v2

@@
Abstract.
@@
In remote run-time attestation, including Control Flow Attestation (CFA) and Data Flow Attestation (DFA), a remote Verifier ($\sf{\mathcal{V}rf}$) requests a potentially compromised Prover device ($\sf{\mathcal{P}rv}$) to generate evidence of its execution control flow path (in CFA) and optionally execution data inputs (in DFA). Recent advances in this space also guarantee that $\sf{\mathcal{V}rf}$ eventually receives run-time evidence from $\sf{\mathcal{P}rv}$, even when $\sf{\mathcal{P}rv}$ is fully compromised. Reliable delivery, in theory, enables secure run-time auditing in addition to attestation, in which $\sf{\mathcal{V}rf}$ examines the evidence to identify exploited vulnerabilities and pinpoint/remediate their root cause(s). However, prior work has left $\sf{\mathcal{V}rf}$’s perspective in this security service unexplored, primarily focusing on methods to securely generate authentic run-time evidence on the (potentially compromised) $\sf{\mathcal{P}rv}$. In this work, we argue that run-time attestation/auditing is only effective if $\sf{\mathcal{V}rf}$ can analyze the received evidence. From this premise, we characterize the different types of evidence produced by run-time attestation/auditing architectures by $\sf{\mathcal{V}rf}$’s ability to use them for vulnerability detection/remediation. As a case study, we propose SABRE– a Security Analysis and Binary Repair Engine – that enables $\sf{\mathcal{V}rf}$ to use run-time evidence to detect control flow attacks, to pinpoint specific instructions that corrupted control data, and to automatically generate binary patches to buffer overflow and use-after-free vulnerabilities without source code knowledge.
1. Introduction
@@
Embedded devices exist at the edge of larger systems, often serving as remotely operated wireless sensors and on-demand actuators. Due to low energy and cost requirements, they are commonly implemented using low-end micro-controller units (MCUs) that lack advanced architectural security features (e.g., virtual memory or memory management units). Consequently, MCUs are more vulnerable and frequently targeted by cyber attacks (Nafees et al., 2023; Kayan et al., 2022).
@@
Given their low budget for preventative security features, MCUs need inexpensive means to prove their software integrity to back-ends that rely on their services (e.g., a device owner or control center). Towards this goal, Remote Attestation (RA(Coker et al., 2011) has been proposed to give a resource-rich back-end – called a Verifier ($\sf{\mathcal{V}rf}$) – the ability to remotely assess the software state of a resource-constrained Prover MCU ($\sf{\mathcal{P}rv}$). In RA, $\sf{\mathcal{V}rf}$ challenges $\sf{\mathcal{P}rv}$ to return a cryptographic proof of $\sf{\mathcal{P}rv}$’s currently installed software binary. Only after receiving an authentic proof will $\sf{\mathcal{V}rf}$ trust that $\sf{\mathcal{P}rv}$ is installed with the correct software.
@@
Although RA can effectively detect illegal code modifications, it cannot detect run-time attacks that do not modify code (Abera et al., 2016). For example, an adversary ($\sf{\mathcal{A}dv}$) could exploit a memory vulnerability (e.g., a buffer overflow (Cowan et al., 2000)) to corrupt control data, including return addresses, indirect jump targets, and function pointers. Consequently, $\sf{\mathcal{A}dv}$ can chain out-of-order instruction sub-sequences to execute (often Turing-complete) malicious behavior (Bletsch et al., 2011; Roemer et al., 2012). Since these attacks do not modify code, they are oblivious to RA.
@@
To address this limitation, RA can be extended into Control Flow Attestation (CFA). In CFA, $\sf{\mathcal{V}rf}$ obtains cryptographic proof of both the installed software image and the exact control flow path followed in the most recent execution of the software (Ammar et al., 2024). This evidence and the respective proof are generated/authenticated by a root of trust (RoT) within $\sf{\mathcal{P}rv}$ that securely records control flow transfer destinations. Therefore, CFA evidence can, in principle, allow $\sf{\mathcal{V}rf}$ to detect both code modification and control-flow hijacking attacks. Recent architectures augment CFA to obtain run-time auditing (Caulfield et al., 2023, 2024a) by guaranteeing the delivery of CFA evidence to $\sf{\mathcal{V}rf}$ even when $\sf{\mathcal{P}rv}$’s software is compromised. See Sec. 2 for details on CFA and run-time auditing RoTs.
@@
We observe that prior work in this area (Ammar et al., 2024) has focused on secure design/implementation of $\sf{\mathcal{P}rv}$ RoTs, often overlooking $\sf{\mathcal{V}rf}$’s role. To our knowledge, no systematic analysis has explored $\sf{\mathcal{V}rf}$’s ability to interpret CFA evidence beyond basic path validation (i.e., determining whether the reported control-flow path is valid). However, in practice, after deeming a path invalid consequently detecting a compromise on $\sf{\mathcal{P}rv}$, $\sf{\mathcal{V}rf}$ would need to use the CFA evidence to (1) identify the root cause of this compromise and (2) implement patches to mitigate it, neither of which has been thoroughly investigated in the literature. These challenges become even more difficulty when $\sf{\mathcal{V}rf}$ does not fully own $\sf{\mathcal{P}rv}$’s software and its executable contains proprietary dependencies/third-party libraries with unavailable source code: a common occurrence in embedded system software chains. Given the prevalence of memory-unsafe languages in embedded software and their various associated exploits ((2023), CISA), we argue that automated methods to identify unknown memory safety vulnerabilities from attested run-time evidence can result in timely detection of otherwise oblivious attacks and elimination of their causes.
While techniques have been proposed for root cause vulnerability analysis in high-end devices (e.g., based on Intel PT traces (Intel, 2015)), none of them can be used to analyze execution traces produced by CFA. In particular, these techniques (Yagemann et al., 2021b, a; Zeng et al., 2022; Liu et al., 2022) require the RoT to provide initial memory snapshots for a remote server to perform a root cause analysis. This requirement is unlikely to be satisfied in the CFA settings due to two reasons. First, this would require $\sf{\mathcal{P}rv}$ to transmit potentially large memory snapshots during CFA, conflicting with the resource-constrained nature of MCUs (Ammar et al., 2024). Second, this snapshot must be stored by $\sf{\mathcal{P}rv}$ during the attested execution. This, in turn, unnecessarily reduces memory available to regular tasks on $\sf{\mathcal{P}rv}$. We expand on these points in Sec. 6. Due to these conflicts, existing root cause analysis frameworks are not directly compatible with CFA.
Our work bridges these gaps with the following contributions:
@@
    We analyze and classify the types of run-time evidence generated by existing CFA RoTs. Their trade-offs are examined in detail with particular focus on their ability to support $\sf{\mathcal{V}rf}$’s remote vulnerability detection and remediation.
@@
    As a case study, we present SABRE: a Security Analysis and Binary Repair Engine based on CFA evidence. SABRE combines binary analysis with CFA run-time evidence to automatically locate and patch buffer overflow and use-after-free vulnerabilities without requiring source-code knowledge.
@@
    Prior root cause analysis assumes $\sf{\mathcal{V}rf}$ has the entire execution context, including an initial memory snapshot. As discussed earlier, this assumption does not hold for CFA. To address this, SABRE takes a different approach: upon detecting corrupted control data (e.g., return addresses or indirect call targets), SABRE performs an additional backward definitions analysis to reconstruct a symbolic memory state leading to the corruption. Since the recovered state is not fully concrete, existing root cause analysis techniques cannot be directly applied. SABRE overcomes this by extending prior work’s forward symbolic data flow analysis to handle symbolic information. Based on this analysis and the detected vulnerability, SABRE automatically generates binary patches and validates their effectiveness.
  4. (4)
    We develop a fully functional open-source prototype of SABRE, supporting TI MSP430 and ARM Cortex-M MCU binaries (Caulfield et al., 2024b) and run-time evidence generated by existing open-source run-time attestation RoTs (based on Trusted Execution Environments (TEEs) (Caulfield et al., 2024a) and custom hardware (Caulfield et al., 2023)).
2. Background
2.1. Remote Attestation
@@
    Upon receiving $Chal$, $\sf{\mathcal{P}rv}$ computes a measurement over its program memory and $Chal$ to produce report $H$.
@@
Early “software-based” attestation (Kennell and Jamieson, 2003; Seshadri et al., 2004, 2005, 2008) methods performed step (2) without relying on secrets and instead incorporated measurable architectural side-effects (e.g., execution/response time) along with the measurement of program memory. To make attestation more suitable to remote wireless connections with unpredictable network latency, subsequent methods required $\sf{\mathcal{P}rv}$ to respond with an authenticated integrity token — e.g., a message authentication code (MAC) or a digital signature – computed over program memory using a secret key (either a pre-shared symmetric key or the private key for a public key $\sf{\mathcal{V}rf}$ knows). Since threat models of RA (and of CFA consequently) assume that $\sf{\mathcal{P}rv}$ is susceptible to full software compromise, the secret key used in this operation must be securely stored by $\sf{\mathcal{P}rv}$’s RoT. Secure storage for the RA secret key implies some level of hardware support for the RoT implementation, ensuring that the key is unmodifiable and inaccessible to any untrusted software running on $\sf{\mathcal{P}rv}$. Therefore, “hardware-based” approaches incorporate dedicated hardware RoT for storing secret keys and computing the measurement (Petroni Jr et al., 2004; Kovah et al., 2012; Schellekens et al., 2008; Vliegen et al., 2019; Noorman et al., 2017). Since such hardware is too costly for low-end MCUs, recent approaches have proposed “hybrid” RoTs (Nunes et al., 2019; Eldefrawy et al., 2012; Brasser et al., 2015), which use hardware for storing secret keys and trusted software for performing the authenticated measurement.
2.2. Code Reuse and Control Flow Attacks
Control flow attacks are made possible due to an exposed memory vulnerability, such as common and pervasive instances like buffer overflow and use-after-free vulnerabilities ((2023), MITRE). Buffer overflows are common in embedded systems software since it is often written in unsafe memory languages like C/C++. For example, recent CVE’s (National Vulnerability Database, 2021b, 2020b, 2020c) have disclosed buffer overflows exploited for arbitrary code execution in IoT devices (Realtek Semiconductor Corp., 2021) and embedded real-time operating systems (Zephyr Project, 2016). Use-after-free is another class of vulnerabilities that can lead to full software compromise. As it is not uncommon for embedded software and RTOSes to use dynamic (i.e., heap) allocation (Masmano et al., 2003; Ramakrishna et al., 2008; Deligiannis and Kornaros, 2016), they can also be vulnerable to use-after-free attacks, which is evidenced by recently disclosed CVE’s (National Vulnerability Database, 2017, 2021a). Since these attacks do not require code modifications, they would remain undetected by classic RA mechanisms.
Verifier ($\sf{\mathcal{V}rf}$) (4) Verify Report (1) Request3) Report Prover ($\sf{\mathcal{P}rv}$) (2) Measure state
Figure 1. RA interaction
2.3. Run-time Attestation and Auditing
@@
Run-time Attestation and Auditing are a set of security services that aim to generate and guarantee the delivery of accurate and authentic evidence of a $\sf{\mathcal{P}rv}$’s run-time behavior to a remote $\sf{\mathcal{V}rf}$, thus reflecting to $\sf{\mathcal{V}rf}$ whether any control flow attack has occurred. One type of run-time attestation is Control Flow Attestation (CFA(Nunes et al., 2021b; Sun et al., 2020; Dessouky et al., 2018; Zeitouni et al., 2017; Toffalini et al., 2019; Zhang et al., 2021; Abera et al., 2016; Geden and Rasmussen, 2019; Huo et al., 2020; Neto and Nunes, 2023; Yadav and Ganapathy, 2023; Armanuzzaman et al., 2025), which extends RA to detect run-time attacks in addition to illegal binary modifications. Upon receiving a request to execute some software function along with a RA challenge from $\sf{\mathcal{V}rf}$, $\sf{\mathcal{P}rv}$ executes the requested function, and a RoT in $\sf{\mathcal{P}rv}$ records a trace of this execution into protected memory. In the most fine-grained case, the trace is a control flow log ($\mathcal{C}F_{Log}$) of all control flow transfers that have occurred during the requested execution. After producing $\mathcal{C}F_{Log}$, $\sf{\mathcal{P}rv}$ performs the authenticated RA integrity measurement over the $\mathcal{C}F_{Log}$, the received challenge, and its own program memory to produce a timely and authenticated report. This report attests to the static state of $\sf{\mathcal{P}rv}$’s program binary and the run-time state within this timely execution. When applicable, data inputs – in Data Flow Attestation (DFA(Sun et al., 2020; Nunes et al., 2021a; Dessouky et al., 2018) – or produced outputs (Nunes et al., 2020; Caulfield et al., 2022) can also be included in this proof to bind obtained results to the proper execution of respective software functions on expected inputs.
@@
Existing CFA techniques use either (1) binary instrumentation along with TEE support; or (2) custom hardware modifications to generate $\mathcal{C}F_{Log}$ by detecting and saving each branch destination to a dedicated and protected memory region. For techniques that use binary instrumentation, a pre-processing phase modifies the binary so that all branch instructions (e.g., jumps, returns, calls, etc.) are prepended with additional calls to a TEE-protected trusted code. Once called, the trusted code appends $\mathcal{C}F_{Log}$ to update the trace with the current branch destination. In hardware-based techniques, custom hardware interfaces with the MCU core to detect branches and records their destinations to a reserved memory region. Recent work (Caulfield et al., 2023, 2024a) has also created mechanisms to implement reliable communication with $\sf{\mathcal{V}rf}$ as a part of the CFA RoT, enabling run-time auditing by ensuring evidence delivery even in the presence of malware that infects $\sf{\mathcal{P}rv}$ and refuses to deliver evidence about its actions/presence.
3. CFA Evidence and its Utility
3.1. Evidence Types